# Magic Mirror on the Wall, How to Benchmark Quantum Error Correction Codes, Overall ?

### v3

@@ V-B Introducing Our Tool for Strategic Recommendation
Our tool primarily aims to match the specific characteristics of a quantum computing scenario with suitable QECCs, recommending a list ranked from most to least favorable. This recommendation also includes the maximum feasible distance for each QECC, aligning with the understanding that longer distances typically correlate with lower logical error rates. Thus, our emphasis on maximizing distance is directly tied to achieving the lowest possible logical error rates. The framework evaluates scenarios based on five key factors: the qubit type being used, the available number of qubits for error correction within the system, the count of qubits in the circuit requiring correction, the presence of multi-qubit gates in the circuit, and the types and rates of physical errors encountered (including Pauli errors like bit-flip, phase-flip, or all-Pauli, as well as depolarizing, gate, and readout errors). Gate errors are the most weighted due to their extensive impact on the quantum system, followed by depolarizing and readout errors. Since all quantum errors can be categorized as Pauli errors, scenarios considering ‘all-Pauli’ errors are deemed most realistic. This flexibility aids in the versatility and characterization of QECCs. Each QECC includes a set of qubit types on which it has been implemented as one of its parameters. The ‘simulation’ type is consistently applied across all QECCs to accommodate those not yet realized in specific scenarios, ensuring that simulated implementations are considered for every QECC. The framework aims to optimize for reduced complexity in overall operations and lattice surgery, while enhancing the code distance and the availability of decoding algorithms. It generates a ranked list of QECCs from most to least suitable based on these considerations, along with the maximum achievable code distance for each QECC, determined by the user’s available qubits for error correction and the qubits in the original circuit. The fundamental operation of the framework is illustrated in Figure 14, providing a foundational overview.
@@
| Parameters | Scenario 1 | Scenario 2 | Scenario 3 | ||||||
| User Input | qType | supercond. | rydberg | simulation | |||||
| maxQAvail | 100 | 600 | 1500 | ||||||
| qOrig | 1 | 2 | 5 | ||||||
| multiQGate? | no | yes | no | ||||||
| errType | bit-flip | all-pauli | phase-flip | ||||||
| depErr | 1E -04 | 1E -04 | 1E -03 | ||||||
| gateErr | 1E -03 | 1E -03 | 1E -02 | ||||||
| readErr | 1E -02 | 1E -02 | 1E -01 | ||||||
| Recommendations |
HeavyH, 11 | ||||||
@@ VI Conclusion
This paper presents the inaugural comprehensive benchmarking framework for Quantum Error Correcting Codes (QECCs), filling a critical void in the field of quantum computing research. Through a detailed comparison of eight QECCs across eight meticulously chosen parameters, our analysis not only highlights the pivotal role of these parameters but also offers an in-depth perspective on the performance of each QECC under different conditions. The innovative use of radar charts for illustrating comparisons clearly demonstrates the unique compromises associated with each QECC, underscoring the intricate process of QECC selection as an exercise in balancing a variety of factors. This study emphasizes the continuous need for benchmarking amidst the discovery of new QECCs and their variants, showcasing the fluid nature of quantum computing where the choice of QECC is highly contextual and no universal solution exists.
Furthermore, we propose a universal, structured methodology for selecting the most appropriate QECC for specific scenarios, aiding users in navigating the complex landscape of quantum computing to find a QECC that aligns with their unique needs. A significant achievement of this research is the creation of a flexible and dynamic pioneering tool that evaluates given scenario characteristics to recommend a prioritized list of QECCs, alongside the maximum achievable distance for each suggested code. This tool is built to accommodate future expansions, including the integration of new QECCs and the easy adjustment of their parameters, ensuring its ongoing utility and adaptability in the face of quantum computing’s rapid advancement. This work is a step forward in ensuring that QECC choices are informed, balanced, and adaptable to evolving quantum computing needs. 222GitHub Repository Link: Magic Mirror

### v4

@@ V-B Strategic Recommendation Tool
Our tool primarily aims to match the specific characteristics of a quantum computing scenario with suitable QECCs, recommending a list ranked from most to least favorable. This recommendation also includes the maximum feasible distance for each QECC, aligning with the understanding that longer distances typically correlate with lower logical error rates. Thus, our emphasis on maximizing distance is directly tied to achieving the lowest possible logical error rates. The framework evaluates scenarios based on five key factors: the qubit type being used, the available number of qubits for error correction within the system, the count of qubits in the circuit requiring correction, the presence of multi-qubit gates in the circuit, and the types and rates of physical errors encountered (including Pauli errors like bit-flip, phase-flip, or all-Pauli, as well as depolarizing, gate, and readout errors). Gate errors are the most weighted due to their extensive impact on the quantum system, followed by depolarizing and readout errors. Since all quantum errors can be categorized as Pauli errors, scenarios considering ‘all-Pauli’ errors are deemed the most realistic. This flexibility aids in the versatility and characterization of QECCs. Each QECC includes a set of qubit types on which it has been implemented as one of its parameters. The ‘simulation’ type is consistently applied across all QECCs to accommodate those not yet realized in specific scenarios, ensuring that simulated implementations are considered for every QECC. The framework aims to optimize for reduced complexity in overall operations and transversal gates while enhancing the code distance and the availability of decoding algorithms. It generates a ranked list of QECCs from most to least suitable based on these considerations, along with the maximum achievable code distance for each QECC, determined by the user’s available qubits for error correction and the qubits in the original circuit.
Parameters | Scenario 1 | Scenario 2 | Scenario 3 | |||||||||
User Input | qType | supercond. | rydberg | simulation | ||||||||
maxQAvail | 100 | 600 | 1500 | |||||||||
qOrig | 1 | 2 | 5 | |||||||||
multiQGate? | no | yes | no | |||||||||
errType | bit-flip | all-pauli | phase-flip | |||||||||
depErr | 1E -04 | 1E -04 | 1E -03 | |||||||||
gateErr | 1E -03 | 1E -03 | 1E -03 | |||||||||
readErr | 1E -02 | 1E -02 | 1E -01 | |||||||||
Recommendations |
@@ VI Conclusion
This study presents a detailed benchmarking framework for Quantum Error Correcting Codes (QECCs), conducting a comparative analysis of nine QECCs across eight essential parameters. Our analysis reveals the significant impact of these parameters on QECC performance under various conditions and the trade-offs involved in selecting a QECC, highlighting the nuanced decision-making required. The study underscores the importance of continuous benchmarking in quantum computing’s evolving landscape, where the choice of QECC is context-dependent. Furthermore, we propose a structured approach for QECC selection tailored to specific needs, offering a flexible and adaptable tool that generates a prioritized list of QECCs based on scenario characteristics and their maximum achievable distance. 222GitHub Repository Link: Magic Mirror
Acknowledgment
The work is supported in parts by the National Science Foundation (NSF) (CNS-1722557, CCF-1718474, OIA-2040667, DGE-1723687, & DGE-1821766) and gifts from Intel. We extend our sincere gratitude to Dr. Anupam Chattopadhyay for his valuable time, wonderful ideas, and insightful suggestions that greatly contributed to this work.